OpenStudy (anonymous):

Evaluate x/y for x= 2/3 and y= 6/7

OpenStudy (anonymous):

7/9 8/10 9/12 12/21

OpenStudy (anonymous):

mine are barly different I had 5/7 not 6/7

OpenStudy (anonymous):

Man :c How do you even solve it?

OpenStudy (anonymous):

go look threw my asked questions should be in their

OpenStudy (anonymous):

Oh wow, that's easy lmao.

OpenStudy (anonymous):

But that would be 18/14 and that's not one of the options.

OpenStudy (shamil98):

when you divide fractions you multiply by the reciprocal

OpenStudy (anonymous):

Yeah, I did that.

OpenStudy (anonymous):

6/7 * 3/2

OpenStudy (shamil98):

\[\huge \frac{ 2 }{ 3 } \div \frac{ 6 }{ 7 }\] \[\huge \frac{ 2 }{ 3 } \times \frac{ 7 }{ 6 }\] what do u get

OpenStudy (anonymous):

14/18

OpenStudy (shamil98):

yes.

OpenStudy (anonymous):

That's not one of the options.

OpenStudy (anonymous):

7/9 8/10 9/12 12/21

OpenStudy (shamil98):

reduce it

OpenStudy (shamil98):

14/18 = 7/9

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Oh, right.